mikewwe13:

Part A There is a function f, where f(n) = 1200n + 3500 represents the number of people in a town who owned a smartphone n years after 2007. How many people owned a smartphone in this town in the year 2010 ? Enter your answer in the box ____________ Part B There is a function p, where p(x) represents the percent of adults in the United States who owned a smartphone x years after 2011. What does p(5) = 68 represent? Select from the drop-down menus to correctly complete the sentence. In the year ___________ 2011 or 2016 ____________ 5 or 68 or 5% or 68% of the adults in the United States owned a smartphone.

Ultrilliam:

@Sam

Vocaloid:

@mikewwe13 n represents years after 2007 so 1 year after 2007, it's 2008 2 years after 2006, it's 2009 3 years after 2006, it's 2010 therefore, when it's 2010, n = 3

Vocaloid:

so let n = 3 and calculate f(n) = 1200n + 3500 = ?

mikewwe13:

100(12n + 35)

Vocaloid:

n = 3 means you can replace "n" with 3 when 3 shows up write this f(n) = 1200*n + 3500 replace "n" with 3 then solve.

mikewwe13:

f(3) = 1200n + 3500 = f = 400n + 3500/3

Vocaloid:

just ignore the f(n) part for now

Vocaloid:

1200*n + 3500 = ?

mikewwe13:

what about the x with n ?

Vocaloid:

there is no x here

Vocaloid:

write this down 1200*n + 3500 = ? replace "n" with 3 then figure out what the expression is equal to

Vocaloid:

* means multiply

mikewwe13:

7100

mikewwe13:

7,100

Vocaloid:

good so that's your answer for part 1

mikewwe13:

7,100

Vocaloid:

Part B There is a function p, where p(x) represents the percent of adults in the United States who owned a smartphone x years after 2011. What does p(5) = 68 represent? Select from the drop-down menus to correctly complete the sentence. In the year ___________ 2011 or 2016 ____________ 5 or 68 or 5% or 68% of the adults in the United States owned a smartphone.

Vocaloid:

p(5) = 68 means that x = 5 and p(5) = 68, which means that the percentage of people who owned a smartphone 5 years after 2011 is 68%

Vocaloid:

what year is it, if it is 5 years after 2011?

mikewwe13:

2006

Vocaloid:

add, don't subtract

Vocaloid:

2011 + 5 = ?

mikewwe13:

2016

mikewwe13:

i meant that at first

Vocaloid:

good so your first blank is 2016 and your second blank is 68%

mikewwe13:

Each day, Ben wants to consume at least 350 grams of protein and carbohydrates combined. He also wants the amount of carbohydrates he consumes to be no more than 3 times the amount of protein. Part B Which combinations allows Ben to consume at least 350 grams of protein and carbohydrates combined while limiting the carbohydrates to no more than 3 times the amount of protein ? Select all that apply. A. 70 grams of protein and 270 grams of carbohydrates B. 75 grams of protein and 275 grams of carbohydrates C. 80 grams of protein and 260 grams of carbohydrates D. 90 grams of protein and 270 grams of carbohydrates E. 100 grams of protein and 290 grams of carbohydrates

Vocaloid:

for each letter choice, add the grams of protein and carbohydrates for example, for A, add 70 and 270 repeat this process for each letter and let me know what you get use a template like this: A = 70 + 270 = 340

mikewwe13:

340

Vocaloid:

good, keep going for B, C, D, E, etc.

mikewwe13:

B. 350 C. 340 D. 360 E. 390

Vocaloid:

good, so if we want the sum of the grams to be ~at least~ 350, we can eliminate one letter. Which letter has a sum less than 350?

mikewwe13:

60

Vocaloid:

which letter, please?

Vocaloid:

A. 340 B. 350 C. 340 D. 360 E. 390 which two letters have a number less than 350?

mikewwe13:

C.

Vocaloid:

there's one more, which one?

mikewwe13:

B.

Vocaloid:

350 is not less than 350, so not B

Vocaloid:

which letters have a number LESS than 350?

mikewwe13:

D.

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid A. 340 B. 350 C. 340 D. 360 E. 390 which two letters have a number less than 350? \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

we are looking for numbers ~SMALLER~ than 350

mikewwe13:

C. D.

Vocaloid:

please try again.

mikewwe13:

C. E.

Vocaloid:

please read what I am writing, slowly and carefully.

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid we are looking for numbers ~SMALLER~ than 350 \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

please slow down

mikewwe13:

A. and D.

Vocaloid:

keep trying.

Vocaloid:

slow down.

mikewwe13:

A. and C.

Vocaloid:

good! A and C are eliminated, so we only need to focus on B, D and E for the next part

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mikewwe13 Each day, Ben wants to consume at least 350 grams of protein and carbohydrates combined. He also wants the amount of carbohydrates he consumes to be no more than 3 times the amount of protein. Part B Which combinations allows Ben to consume at least 350 grams of protein and carbohydrates combined while limiting the carbohydrates to no more than 3 times the amount of protein ? Select all that apply. A. 70 grams of protein and 270 grams of carbohydrates B. 75 grams of protein and 275 grams of carbohydrates C. 80 grams of protein and 260 grams of carbohydrates D. 90 grams of protein and 270 grams of carbohydrates E. 100 grams of protein and 290 grams of carbohydrates \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

now, we want the grams of carbohydrates to be less than 3 times the grams of protein for B, D, and E, calculate: 3 * grams of protein

mikewwe13:

B. 825 D. 810 E. 870

Vocaloid:

3 * grams of ~protein~ not carbohydrates please

mikewwe13:

oh wait i did it wrong

mikewwe13:

B. 225 D. 240 E. 300

Vocaloid:

check your numbers again

Vocaloid:

B has 75 g protein D has 90 g protein E has 100 g protein

Vocaloid:

multiply these by 3

mikewwe13:

3 x 75 right ?

Vocaloid:

just check D again, C and E were right

mikewwe13:

3 X 100 ?

mikewwe13:

90

Vocaloid:

good, and 90*3 = ?

mikewwe13:

270

Vocaloid:

good so B = 225g D = 270g E = 300g so we want these numbers to be less than the carbohydrate numbers

Vocaloid:

*sorry, greater than

Vocaloid:

now, for B, there are 275g carbs D = 270g carbs E = 290 g carbs let's compare them side by side

Vocaloid:

B: 225g protein vs 275g carbs is protein or carbs bigger?

mikewwe13:

carbs is bigger

Vocaloid:

good, so B cannot be one of the choices

Vocaloid:

let's check D 270g protein vs 270g carbs they're equal, so D works as an answer lets check E

Vocaloid:

300g protein vs 270 carbs is protein or carbs bigger?

mikewwe13:

protein

Vocaloid:

good so D and E are the answers

Vocaloid:

can we open a new question please

mikewwe13:

ok